Find and classify the critical points off(x)=9x4(2−x)5 as local maxima and minima.
Critical points: x
1
Expert's answer
2020-03-04T17:02:52-0500

The critical points at f'(x)=0,

"(9*x^4)'*(2-x)^5+9*x^4*((2-x)^5)'=0,\n\n36*x^3*(2-x)^5-45*x^4*(2-x)^4=0, \n\n9*x^3*(2-x)^4*(8-9*x)=0,"

"f'(-1)<0; \n\n\nf'(0,5)>0;\n\n\nf'(1)<0;\n\n\nf'(3)<0."

When passing through point 0, the derivative changes the sign from - to +, then

x1=0 is a point of minimum.

When passing through point 8/9, the derivative changes the sign from + to -, then

x2=8/9 is a point of maximum.

When passing through point 0, the derivative does not change the sign, then

x3=2 is not a point of minimum/maximum.